# Explainable Concept Generation through Vision-Language Preference Learning

### v2

@@ 3.2 Extracting Seed Prompts
Since a generative model can generate arbitrary images, if we provide good starting point for optimization then the convergence to explainable states would be faster. In this paper, to extract seed prompts for a particular class we make use of the off-the-shelf VQA model followed by several preprocessing steps, as described in Appendix C.3. We also explore how random gibberish prompts can be used as seed prompts in Appendix C.4 which did not yield useful concepts.
@@ 3.3 Deep Reinforcement Learning Formulation
Our objective of using deep RL is automatically controlling text input of Stable Diffusion. As text input, we start with seed prompts from Section 3.2, $\mathcal{K}$, that have the potential to generate meaningful concept images after many deep RL episodes. We setup our RL state-action at iteration $t$ as,
@@
  • Reward $r_{t}$: Reward $r_{t}$ is proportional to the TCAV score computed at state $s_{t}$ on action $a_{t}$, adjusted by a monotonically increasing scaling factor $\xi_{t,k}$. As each seed concept reaches the explainable state at different times, this factor is introduced to scale the reward over time $t$ for each unique seed concept $k$. Since the $g(.)$ is getting optimized at each time step $t$. The scaling factor is updated as $\xi_{t+1,k}\leftarrow\min(1,\frac{\xi_{t,k}+1}{T})$, where T is total number of RL steps. Therefore, the expected cumulative adjusted reward is $R(\pi)=\mathbb{E}[\sum_{t=0}^{T}\xi_{t}\cdot r_{t}(s_{t},a_{t})]$.
@@
$Q^{*}(s,a)=\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[\xi_{t}r(s,a)+\gamma\max_{a^{\prime}\in A}Q_{\text{target}}(s^{\prime},a^{\prime})].$ (3)
@@ 3.4 Optimizing the States
At time $t$, the policy picks the seed prompt $k_{t}$, which is then used by the generative model, $g(k_{t};w_{t})$, with model weights $w$, to generate $2Z$ number of images. We randomly divide the generated images into two groups: $X_{c_{1},t}=\{x_{c_{1},t,i}\}_{i=1}^{Z}$ and $X_{c_{2},t}=\{x_{c_{2},t,i}\}_{i=1}^{Z}$. Let the TCAV scores of each group be $TS_{c_{1},m,t}$ and $TS_{c_{2},m,t}$. Since our objective is to find concepts that generate a higher TCAV score, concept images that have a higher score is preferred. Note that, unlike in the classical preference optimization setting with a human to rank, RLPO preference comes from the TCAV scores (e.g., $TX_{c_{1},t}\succ TX_{c_{2},t}$). We call this notion RLPO-XAI in some ablation studies below. If the generative model at time $t$ is not capable of generating concepts that are in an explainable state, $\max(TS_{c_{1},m,t},TS_{c_{2},m,t})\leq\eta$, we then perform preference update on SD’s weights (more details in Appendix C.5). Following Low-Rank Adaptation (LoRA) (Hu et al., 2021), we only learn auxiliary weights $a$ and $b$ at each time step, and update the weights as $w_{t+1}\leftarrow w_{t}+\lambda ab$.
As the deep RL agent progresses over time, the states become more relevant as it approaches explainable states (Fig. 1), thus the same action yields increasing rewards over time. To accommodate this, with reference to the rewards defined in Section 3.3, we introduce a parameter $\xi$, which starts at 0.1 and incrementally rises up to 1 as the preference threshold, $\eta$, is approached. Different actions may result in different explainable states, reflecting various high-level concepts inherent to $f(\cdot)$. Some actions might take longer to reach an explainable state. Also, it is possible for different actions to lead to the same explainable state. As the goal is to optimize all states to achieve a common target, DQN progressively improves action selection to expedite reaching these states. Thus, deep RL becomes relevant as it optimizes over time to choose the actions that are most likely to reach an explainable state more efficiently.
@@ 4 Experiments
To test the effectiveness of our approach, we tested it across multiple models and several classes. We considered three CNN-based classifiers, ResNet-50 (He et al., 2016), GoogleNet (Szegedy et al., 2015), and InceptionV3 (Szegedy et al., 2016), and two transformer-based classifiers, ViT (Dosovitskiy et al., 2020) and Swin (Liu et al., 2021), pre-trained on ImageNet dataset. Unless said otherwise, only GoogleNet results are shown in the main paper. All other model details and results are provided in Appendix C.1 and D.3), respectively. All metrics are defined in Appendix B.1.
4.1 Ablation study: Search Strategies (Why deep RL?)
We chose DQN as our RL algorithm because of its ability to effectiveness traverse through discrete action space Mnih et al. (2015) (20 unique seed prompts). We assess the effectiveness of RL by disabling the preference optimization step. As shown in Table 2, on GoogleNet classifier, compared to $\epsilon$-greedy methods, RL setup exhibits a higher entropy, average normalized count (ANC), and inverse coefficient of variance (ICV) (See Appendix B.1 for definitions), indicating RL’s ability to take diverse actions that results in diverse concepts.
4.2 Ablation study: Scoring Mechanisms (Why TCAV?)
Another important aspect of our setup is the use of TCAV score, an XAI method, to provide preference feedback and calculate rewards. Alternatively, this XAI scoring feedback can also be replaced with human feedback or LLM-based AI feedback. As an additional experiment, to test the effect of human feedback, we conducted human feedback experiment with eight human subjects who provided live human feedback. Further, to evaluate the LLM-based AI feedback, we made use of GPT-4o. More details on the experiment setup and results are provided in Appendix D.2. As shown in Table 4, we concluded that, even though other feedback techniques can be used, XAI-based feedback is best for generating concepts that are important to model with high speed and low computation cost. Though human and AI (GPT4o) are good at correlating semantics, by only looking at test images and concepts instead of model activations, they are not able to provide model specific explanations. Also, human experiments tend to be expensive.
Table 1: Search strategy ablation. We see that RL, compared to $\epsilon$-greedy search, is the best strategy to efficiently explore the search space with high entropy, high average normalized count (ANC) per action, and high inverse coefficient of variance (ICV).
@@ 4.2 Ablation study: Scoring Mechanisms (Why TCAV?)
Table 2: Exploration Gap (EG) and Odds calculated based on the responses for ours and retrieval based method, respectively, from the human survey (Appendix D.6).
@@
Method Class-based Model-specific Feedback Execution
Explanations Explanations Cost* Time (↓)
RLPO-HF NIL 180 $\pm$ 30s
RLPO-AIF >10 GB 72 $\pm$ 1.2s
RLPO-XAIF <1 GB 56 $\pm$ 0.7s
* Feedback cost refers to the storage or memory requirements for feedback processing.
Table 4: Novel concepts: $TS_{c,m}$ (TCAV score), CS (Cosine similarity), ED (Euclidean distance), RCS, and RED (CS and ED with ResNet50 embedding)
Model Concepts $TS_{c,m}(\uparrow)$ CS ($\downarrow$) ED ($\uparrow$) RCS ($\downarrow$) RED ($\uparrow$)
@@
4.3 What kind of concepts can RLPO generate?
@@
Novel concepts. As illustrated in Fig. 4, we observed that the RLPO can generate concepts that a human would not typically think of but leads activations of the DNN to trigger. To validate this hypothesis, we conducted a survey to see if humans can think of these generated concepts as important for the neural network to understand a certain class (Table 2, The Exploration Gap (EG) quantifies the proportion of missed optimal actions, defined as $1-\text{Accuracy}$, highlighting how humans frequently miss the most optimal actions when presented with generated concepts.). As detailed in Appendix D.6, we presented a random class image followed by two concepts, one generated by our method and another from a previous retrieval based method (Fel et al., 2024; 2023). We presented choices from the concepts with high score from both methods and we discovered that while most participants could recognize retrieval-based concepts, only those with domain-specific knowledge could identify generated concepts. This indicates that most people can only identify concepts from a small subset of what $f(.)$ learns during training. Intuitively, when we retrieve concepts from the test class, they tend to be similar to the test images.
@@
Figure 4: Samples of concepts (with different TCAV scores, not shown here) generated by different methods. Observe that RLPO generates diverse images, not just patches from the test images.
@@
We also compare the generated concepts qualitatively and quantitatively. Fig. 4 shows the diversity of concepts generated by our method and other retrieval based methods. Additionally, we also verify this by computing the vector similarity of the CLIP and ResNET50 embeddings between $X_{c}$ and $X_{m}$ for multiple retrieval based methods. As highlighted in Table 4, we observe that retrieval based methods tend to have high cosine similarity between extracted concepts and test images, making them less useful as abstract concepts (e.g., to explain the zebra class, a patch of zebra as a concept is less useful compared to stripes concept).
Abstract concepts. In Fig. 5, we observe the progression of output concepts generated by the SD when RLPO is applied for the seed prompt “zoo,” of tiger class. These abstractions hint us about what the model prefers when it is looking for tiger, starting from a four-legged orange furred animal, to black and white stripes with orange furred animal, to black and white stripes with orange furred and whiskers. We obtain concepts with various abstractions by changing $\eta$ (Currently, it is not possible for our method to decide $\eta$ to get a particular level of abstraction).
Figure 5: Different levels of abstractions for the “Tiger” class on GoogleNet classifier. The generated image begins as a random zoo image and gradually transition to images with tiger features. We also see that the model’s prediction changes over time, starting from a random classification of oxcart to classifying the generated concept as tiger itself.
@@
Multiple concepts. Because RLPO algorithm explores various explainable states, we can obtain multiples concepts with varying level of importance. Fig. 6 shows the top three class-level concepts identified by our method for the “zebra” class for the GoogleNet classifier. We see that, each concept set has a different TCAV score associated with them indicating their importance.
@@
Figure 6: The figure shows the concepts identified by our method and where they are located in the input class image (“zebra” class) for GoogleNet classifier. As highlighted the “stripes” concept image are located near zebra, the “running” concept images, showing trees are highlighting in background of the input image, and the “mud” concept highlighting the grass and soil in the input image. The concepts are ordered in their importance (TCAV score) with “stripes” being the highest and “mud” being the lowest for the selected class.
4.4 Are the generated concepts reliable?
@@
After generating the concepts, next step is to identify what those concepts signify. To locate where in the class images generated concepts correspond, we made use of CLIPSeg (Lüddecke & Ecker, 2022), a transformer-based segmentation model which takes in concept images as prompts, $X_{c}$, and highlights in a test image, $x\in X_{m}$, which part resembles the input prompt as a heat map. More details on this is available in Appendix D.3.3. As shown in Fig. 6, class image on left highlights the top 3 identified concepts by RLPO. We also compare the output generated by other popular XAI techniques such as LIME and GradCam with ones generated by RLPO. As shown in Fig. 7, we can see that other methods just explains where the model is looking at whereas our approach also explains what type of features is the model focuses on.
After finding the relationship between generated concepts and input images, we need to validate the importance of the identified concepts. To that end, we applied c-deletion, a commonly used validation method in XAI, to the class images for each identified concept. We gradually deleted concept segments based on the segmentation heat map obtained from ClipSeg. The results for the c-deletion are shown in the Fig. 9. We see the area under curve is the highest for the most important concept “stripes” and the lowest the least important concept “mud,” indicating the order of importance of each concept. More examples on the c-deletion are in Appendix D.3.4.
@@
Figure 8: C-deletion. Removing concepts over time to measure the reliability. The colored numbers indicate the area under the curve (the lower the better as it indicates important concepts are removed sooner.)
Figure 9: How some eight concepts are shifted from pre (blue) to post (orange) fine-tuning of ResNet50 model on Blue Jay class.
Figure 10: We use sentiment analysis in NLP to show the generalizability of RLPO. Here, concepts tend to be synonyms and the numbers indicate the TCAV scores. Generated concepts explain why a given text is classified as a positive sentiment.
4.5 Applications and generalizability
We show how RLPO can be used as a diagnostic tool for the engineers. As a specific application, we see what concepts are removed and added, as well as how the concept importance changes when we fine-tune ResNet50 model on ImageNet to improve accuracy (Fig. 9). More details about the experiment is present in Appendix D.5.

### v3

@@ 3.2 Extracting Seed Prompts
Since a generative model can generate arbitrary images, if we provide good starting point for optimization then the convergence to explainable states would be faster. In this paper, to extract seed prompts for a particular class we make use of the off-the-shelf VQA model followed by several pre-processing steps, as described in Appendix B.3. We also explore how gibberish prompts can be used as seed prompts in Appendix B.4 which did not yield useful concepts.
@@ 3.3 Deep Reinforcement Learning Formulation
Our objective of using deep RL is automatically controlling text input of Stable Diffusion. As text input, we start with $\mathcal{K}$ seed prompts from Section 3.2, that have the potential to generate meaningful concept images after many deep RL episodes. We setup our RL state-action at iteration $t$ as,
@@
  • Reward $r_{t}$: It is proportional to the TCAV score computed at state $s_{t}$ on action $a_{t}$, adjusted by a monotonically increasing scaling factor $\xi_{t,k}$. As each seed concept reaches the explainable state at different times, this factor is introduced to scale the reward over time $t$ for each unique seed concept $k$. Since the $g(.)$ is getting optimized at each time step $t$. The scaling factor is updated as $\xi_{t+1,k}\leftarrow\min\Big{(}1,\frac{\xi_{t,k}+1}{T}\Big{)}$, where T is total number of RL steps. Therefore, the expected cumulative adjusted reward is $R(\pi)=\mathbb{E}\Big{[}\sum_{t=0}^{T}\xi_{t}\cdot r_{t}(s_{t},a_{t})\Big{]}$.
@@
$Q^{*}(s,a)=\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[\xi_{t}r(s,a)+\gamma\max_{a^{\prime}\in A}Q_{\text{target}}(s^{\prime},a^{\prime})]$ (3)
@@ 3.4 Optimizing the States
At time $t$, the policy picks the seed prompt $k_{t}$, which is then used by the generative model, $g(k_{t};w_{t})$, with model weights $w$, to generate $2Z$ number of images. We randomly divide the generated images into two groups: $X_{c_{1},t}=\{x_{c_{1},t,i}\}_{i=1}^{Z}$ and $X_{c_{2},t}=\{x_{c_{2},t,i}\}_{i=1}^{Z}$. Let the TCAV scores of each group be $TS_{c_{1},m,t}$ and $TS_{c_{2},m,t}$. Since our objective is to find concepts that generate a higher TCAV score, concept images that have a higher score is preferred. Note that, unlike in the classical preference optimization setting with a human to rank, RLPO preference comes from the TCAV scores (e.g., $TX_{c_{1},t}\succ TX_{c_{2},t}$). We call this notion RLPO-XAIF in ablation studies below. If the generative model at time $t$ is not capable of generating concepts that are in an explainable state, $\max(TS_{c_{1},m,t},TS_{c_{2},m,t})\leq\eta$, we then perform preference update on SD’s weights (more details in Appendix B.5). Following Low-Rank Adaptation (LoRA) (Hu et al., 2022)—a method that allows quick SD adaptation with a few samples,— we only learn auxiliary weights $a$ and $b$ at each time step, and update the weights as $w_{t+1}\leftarrow w_{t}+\lambda ab$.
As the deep RL agent progresses over time, the states become more relevant as it approaches explainable states (Fig. 2), thus the same action yields increasing rewards over time. To accommodate this, with reference to the rewards defined in Section 3.3, we introduce a parameter $\xi$, which starts at 0.1 and incrementally rises up to 1 as the preference threshold, $\eta$, is approached. Different actions may result in different explainable states, reflecting various high-level concepts inherent to $f(\cdot)$. Some actions might take longer to reach an explainable state. As the goal is to optimize all states to achieve a common target, DQN progressively improves action selection to expedite reaching these states. Thus, deep RL becomes relevant as it optimizes over time to choose the actions that are most likely to reach an explainable state more efficiently.
@@ 4 Experiments
To verify the effectiveness of our approach, we tested it across multiple models and several classes. We considered two CNN-based classifiers, GoogleNet (Szegedy et al., 2015) and InceptionV3 (Szegedy et al., 2016), and two transformer-based classifiers, ViT (Dosovitskiy et al., 2020) and Swin (Liu et al., 2021), pre-trained on ImageNet dataset. Unless said otherwise, only GoogleNet results are shown in the main paper. All other model details and results are provided in Appendix B.1 and C.3, respectively.
4.1 Ablation: Search Strategies (Why Deep RL?)
We chose DQN as our RL algorithm because of its ability to effectiveness traverse through discrete action space (Mnih et al., 2015) (20 unique seed prompts). We assess the effectiveness of RL by disabling the preference optimization step. As shown in Table 1, on the GoogleNet classifier, compared to $\epsilon$-greedy methods, the RL setup exhibits higher entropy, average normalized count (ANC), and inverse coefficient of variance (ICV) (See Appendix A.1 for definitions), indicating RL’s ability to efficiently explore across diverse actions. Qualitative results obtained on updating SD model with and without RL are discussed in Appendix C.5.
4.2 Ablation: Scoring Mechanisms (Why TCAV?)
Another important aspect of our setup is the use of TCAV score, an XAI method, to provide preference feedback and calculate rewards. Alternatively, this XAI scoring feedback can also be replaced with human feedback or LLM-based AI feedback. As an ablation study, to test the effect of human feedback, we conducted human feedback experiment with eight human subjects who provided live human feedback. Further, to evaluate the LLM-based AI feedback, we made use of GPT-4o. More details on the experiment setup and results are provided in Appendix C.2. As shown in Table 2, we concluded that, even though other feedback techniques can be used, XAI-based feedback is best for generating concepts that are important to model with high speed and low computation cost. Though human and AI (GPT4o) are good at correlating semantics, by only looking at test images and concepts instead of model activations, they are not able to provide model specific explanations.
Table 1: Search strategy ablation. We see that RL, compared to $\epsilon$-greedy search, is the best strategy to explore the search space with high entropy, average normalized count (ANC) per action, and inverse coefficient of variance (ICV).
@@ 4.2 Ablation: Scoring Mechanisms (Why TCAV?)
Method Class-based Model-specific Feedback Execution
Explanations Explanations Cost* Time (↓)
RLPO-HF NIL 180 $\pm$ 30s
RLPO-AIF $>10$ GB 72 $\pm$ 1.2s
RLPO-XAIF $<1$ GB 56 $\pm$ 0.7s
* Feedback cost refers to the GPU memory requirements.
Table 3: Exploration Gap (EG) and Odds for different methods based on the human survey (Appendix C.9). This verifies that RLPO can generate concepts that human cannot think of.
@@
Table 4: Novel concepts: $TS_{c,m}$ (TCAV score), CS (Cosine similarity), ED (Euclidean distance), RCS, and RED (CS and ED with ResNet50 embedding). This indicates that RLPO can generate a diverse set of concepts that triggers the network.
Methods Concepts $TS_{c,m}(\uparrow)$ CS ($\downarrow$) ED ($\uparrow$) RCS ($\downarrow$) RED ($\uparrow$)
@@
4.3 Concepts Generated by RLPO
Objective of RLPO is to automatically generate concepts that a human or a retrieval method cannot propose but the neural network has indeed learned (i.e., gets activated). As illustrated in Fig. 4, we observed that the RLPO can generate diverse set of concepts that a human would not typically think of but leads activations of the DNN to trigger. To validate this hypothesis, we conducted a survey to see if humans can think of these generated concepts as important for the DNN to understand a certain class.
@@ 4.3 Concepts Generated by RLPO
As detailed in Appendix C.9, during our human survey, we presented a random class image followed by two concepts, one generated by our method and another from a previous retrieval-based method (Fel et al., 2024, 2023). While the choices were similar in terms of XAI-score from both methods, we discovered that most participants recognize retrieval-based concepts, and only those with domain-specific knowledge could identify generated concepts as important. As highlighted in Table 3, high Exploration Gap (EG) (definition in Appendix A.1) of our method indicates that most people can only identify concepts from a small subset of what $f(\cdot)$ learns during training. Intuitively, when we retrieve concepts from the test class, they tend to be similar to the test images. We further verify that the generated concepts have the following properties.
@@
Figure 4: Samples of concepts generated by different methods. Observe that RLPO generates diverse images, beyond patches from the test images.
@@
Diverse representations per concept. We verify the diversity (e.g., different types of stripes) of generated and retrieval-based concepts by computing the vector similarity of the CLIP and ResNET50 embeddings between $X_{c}$ and $X_{m}$ for different classes. As highlighted in Table 4, we observe that concepts from retrieval-based methods tend to have high cosine similarity with test images, making them less useful as abstract concepts (e.g., to explain the zebra class, a patch of zebra as a concept is less useful compared to stripes concept).
@@
Multiple concepts per class. Since RLPO algorithm explores various explainable states, we can obtain multiples concepts (e.g., stripes, savanna) with varying level of importance. Fig. 5 shows the top three class-level concepts identified by our method for the “zebra” class for the GoogleNet classifier. We see that, each concept set has a different TCAV score associated with them indicating their importance. Additionally, as shown in Table 5, these concepts are inherently different from one another. As an additional result, we can see in Appendix C.4 how each concept evolves—representing different levels of concept abstractions—with each RL step.
Table 5: Inter-concept Comparison for “zebra” class across three trial runs. Low Cosine Similarity (CS), high Wasserstein Distance (WD), and high Hotelling’s T-squared Score (HTS) suggests that concepts generated from different seed prompts are different from one another.
Metrics Stripes-Running Running-Mud Mud-Stripes
Concept Concept Concept
Avg. CS ()($\downarrow$)( ↓ ) $0.67\pm 0.01$ $0.69\pm 0.0004$ $0.73\pm 0.0004$
Avg. WD ()($\uparrow$)( ↑ ) $8.15\pm 0.05$ $7.85\pm 0.02$ $7.48\pm 0.03$
Avg. HTS ()($\uparrow$)( ↑ ) $7598.50\pm 84.5$ $13069.68\pm 2147.81$ $7615.73\pm 538.06$
Are they from the
same distribution? No No No
@@
Figure 5: The figure shows the concepts generated by our method and where they are located in the input image (“zebra” class) for GoogleNet classifier. As highlighted the “stripes” concept images are located near zebra, the “running” concept images, showing trees, highlight the background, and the “mud” concept highlights the grass and soil in the input image. The concepts are ordered in their importance (TCAV score) with “stripes” being the highest and “mud” being the lowest.
4.4 Are Generated Concepts Correct?
@@
After generating the concepts, next step is to identify what those concepts signify. To locate where in the class images generated concepts correspond, we made use of CLIPSeg (Lüddecke & Ecker, 2022), a transformer-based segmentation model which takes in concept images as prompts, $X_{c}$, and highlights in a test image, $x\in X_{m}$, which part resembles the input prompt as a heat map. More details on this is available in Appendix C.3.3. As shown in Fig. 5, class image on left highlights the top 3 identified concepts by RLPO. We also compare the output generated by other popular XAI techniques such as LIME and GradCam with ones generated by RLPO, more details in Appendix C.6.
Table 6: Intra-concept Comparisons for “zebra” class across three trial runs. High Cosine Similarity (CS), low Wasserstein Distance (WD), and low Hotelling’s T-squared Score (HTS) suggests that concepts generated from same seed prompts lies on the same distribution.
Metrics Stripes-Stripes Running-Running Mud-Mud
Concept Concept Concept
Avg. CS ()($\uparrow$)( ↑ ) $0.99\pm 0.0008$ $0.99\pm 0.0004$ $0.99\pm 0.0004$
Avg. WD ()($\downarrow$)( ↓ ) $0.95\pm 0.07$ $0.82\pm 0.07$ $0.82\pm 0.06$
Avg. HTS ()($\downarrow$)( ↓ ) $2.46\pm 0.091$ $2.36\pm 0.08$ $2.66\pm 0.22$
Are they from the
same distribution? Yes Yes Yes
@@
Figure 6: C-deletion. Removing concepts over time to measure the reliability. The colored numbers indicate the area under the curve.
After finding the relationship between generated concepts and input images, we need to validate the importance of the identified concepts. To that end, we applied c-deletion, a commonly used validation method in XAI, to the class images for each identified concept. We gradually deleted concept segments based on the heat map obtained from ClipSeg. The results for the c-deletion are shown in the Fig. 6. We see the area under curve is the highest for the most important concept “stripes” and the lowest for least important concept “mud,” indicating the order of importance of each concept. More examples on the c-deletion are in Appendix C.3.4. Additionally, to verify if the generated concepts are consistent, we compared their CLIP embedding across multiple runs. As shown in Table 6, we can see that the concepts generated from same seed prompt belongs to the same distribution across multiple runs.
4.5 How Are Generated Concepts Useful to Engineers?
To verify the usability of the generated concepts, we conducted a human study with 19 ML engineers. We first provided them the concept generated by our method for “zebra” class and ask them to choose relevant concepts for GoogleNet to classify a zebra without telling them that all shown images are actual concepts. All the engineers selected the “stripes” concept to be important while some also selected the “mud” concept. But most missed the “running” concept. This indicates that engineers cannot think of all the important concepts that gets the neural network activated. In the next step, we showed engineers the concept-explanation mapping on a random input image (similar to Fig. 5) and asked them if the provided explanation helped them understand the model better and if it provided new insights. 94.7% of the engineers agreed that the explanation helped in better understanding the neural network and 84.2% agreed that it provided new insights. This result shows that the new concepts discovered by our proposed method help engineers discover new patterns that they did not imagine before (More details in Appendix C.10).
We now demonstrate how engineers can use these newly revealed information about concepts to improve the model. As identified by RLPO, for Tiger class, the base GoogleNet model gives equal importance to both foreground (highlighted by concepts “orange black and white” and “orange and black”) and background (highlighted by concepts “blurry”) in the input (see Fig. 12 in Appendix C.3 for example explanations). As shown in Fig. 7, when we fine-tune the GoogleNet on images of Tiger-related concepts, we see that the fine-tuned model now focuses more on tiger than the background while maintaining a similar accuracy (65.6%). Details in Appendix C.8.
Figure 7: Usefulness of RLPO. Fine-tuning GoogleNet based on generated concepts for the Tiger class.
4.6 Can concept Generation Help in Understanding Model Bias?
Apart from generating actionable concepts, our proposed framework can also be used to identify spurious correlations or undesirable biases learned by the model during training. To demonstrate this, we applied RLPO to a ResNet-18 classifer trained on the CelebA dataset (Liu et al., 2015) for “Blonde” versus “Not Blonde” classification, a task with known gender biases (De Coninck et al., 2024; Subramanyam et al., 2024; Chen et al., 2023). We adapted our seed prompts to include higher-level semantic about gender, and as shown in Fig. 8, RLPO showed that concepts representing “female face” were more important for the model for the “Blonde” class. Additionally, we see that for the “male face” seed prompt, the generated concepts started generating males with long blonde hair, further narrowing down on the fine-grain concepts the model is looking at (i.e. long and blonde hair).
Figure 8: (a) TCAV scores obtained from “Male face” and “Female face” seed prompt, highlighting the importance of female gender over male. (b) Sample generated concepts for “Female face” and “Male face” seed prompt.
4.7 Can RLPO Generalized Beyond Images?